HIGH RESOLUTION SPECTRUM OF THE ^{13}C^{12}C^{12}C LOWEST BENDING MODE

C. P. ENDRES, V. LUTTER, J. KÖTTING, J. KRIEG, S. THORWIRTH, S. SCHLEMMER, \underline{T. F. GIESEN}, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; M. E. HARD-ING, Karlsruher Institut für Technologie, Institut für Nanotechnologie, 76021 Karlsruhe, Germany; J. VÁZQUEZ, Department of Chemistry and Biochemistry, The University of Texas at Austin, Austin, Texas 78712, USA.

Linear C_{3} is a floppy molecule which possesses an extremely low lying bending mode, ν_{2}, at roughly 60 cm^{−1} or 1.9 THz. Based on highly accurate laboratory data^{a}^{,}^{b}^{,}^{c} C_{3} has been detected in various astronomical sources^{d}^{,}^{e} most recently with the HIFI instrument aboard the Herschel satellite^{f}. Although C_{3} turns out to be quite abundant in interstellar environments which makes a search for ^{13}C substituted isotopologs feasible, other isotopologs could not be detected so far, because no accurate transition frequencies have been available for these species in this frequency range. Relative abundance ratios of C_{3} isotopologs might give important hints on its building mechanism and further constraints for chemical networks. In this work, the spectrum of the ν_{2} lowest bending mode of ^{13}CCC has been investigated. We used laser ablation of ^{13}C enriched carbon samples to record absorption spectra in a supersonic jet expansion. The radiation in our setup is generated by a synthesizer referenced to a Rubidium standard in combination with a frequency multiplier chain and detected by a liquid Helium cooled InSb bolometer. The laboratory search has been supported by high-level coupled-cluster calculations, which turns out to compare very favorably with obtained experimental molecular parameters.

aSchmuttenmaer, C. A., Cohen, R. C., Pugliano, N., Heath, et al., Sience 249, 897−900 (1990)

bGiesen, T. F., van Orden, A. O., Cruzan, J. D., and Provencal, R. A., et al., Astrophys. J. 551, L181−L184 (2001)

cGendriesch, R. and Pehl, K. and Giesen, T. and Winnewisser, G. and Lewen, F., Z. Naturforsch. 58a, 129−138 (2003)

dVan Orden, A., Cruzan, J. D., Provencal, R. A., et al. in Proc. Airborne Astronomy Symp., ASP Conf. Ser. 73, 67 (1995)

eCernicharo, J. and Goicoechea, J. R. and Caux, E., Astrophys. J. Lett. 534, L199−L202 (2000)

fMookerjea, B., Giesen, T., Stutzki, J., Cernicharo, J., et al., Astron. Astrophys. 521, L13 (2010)